# Automated Essay Scoring Incorporating Annotations from Automated Feedback Systems

### v1

@@ Abstract
This study illustrates how incorporating feedback-oriented annotations into the scoring pipeline can enhance the accuracy of automated essay scoring (AES). This approach is demonstrated with the Persuasive Essays for Rating, Selecting, and Understanding Argumentative and Discourse Elements (PERSUADE) corpus. We integrate two types of feedback-driven annotations: those that identify spelling and grammatical errors, and those that highlight argumentative components. To illustrate how this method could be applied in real-world scenarios, we employ two LLMs to generate annotations – a generative language model used for spell-correction and an encoder-based token-classifier trained to identify and mark argumentative elements. By incorporating annotations into the scoring process, we demonstrate improvements in performance using encoder-based large language models fine-tuned as classifiers.
@@ 1 Introduction
Automated Essay Scoring (AES) uses statistical models to assign grades to essays that approximates hand-scoring Shermis and Hamner (2013). Automated Writing Evaluation (AWE) is the provision of automated feedback designed to help students iteratively improve their essays Huawei and Aryadoust (2023). Initial attempts at AES and AWE were based on Bag-of-Words (BoW) models that combine frequency-based and hand-crafted features Attali and Burstein (2006); Page (2003). Well-designed features can serve two purposes; to improve scoring accuracy and provide feedback to students to improve their essays. These features tend to be global features, such as the number of words, sentence length, or readability metrics, and not based on fine-grained semantics or the organizational structure of essays.
Many modern AES engines employ transformer-based Large Language Models (LLM)s Rodriguez et al. (2019), which offer improved accuracy over bag-of-words models. However, this comes at the expense of reduced interpretability due to implicit feature definition. Some researchers have sought to combine LLM-derived features with traditional hand-crafted features to enhance accuracy and provide some level of interpretability Uto and Uchida (2020). LLMs also offer the ability to provide semantically rich feedback such as key phrases from explainable AI Boulanger and Kumar (2020), annotation schemas for automated writing evaluation systems Crossley et al. (2022); Lottridge et al. (2024), and the generation of detailed feedback through LLM prompting techniques Lee et al. (2024). This study considers how these semantically rich features, designed primarily for providing feedback, can also enhance scoring accuracy.
@@
The two classes of features we consider are derived from Grammatical Error Correction (GEC) and Computational Argumentation. The goal of GEC is to provide a mapping from a sentence that may or may not contain errors in language, to a version with the same meaning with fewer errors in language Martynov et al. (2023), while computational argumentation seeks to isolate and analyze the set of argumentative components of an essay Stab and Gurevych (2014a). For these features to be incorporated into an AES pipeline, we leverage the ability of language models to accurately annotate argumentative components Ormerod et al. (2023) and correct spelling and grammatical errors Rothe et al. (2021). These spelling and grammatical corrections can then be annotated and classified to provide locally defined information on conventions Korre and Pavlopoulos (2020). Once these annotations have been derived, we incorporate the annotations using Extensible Markup Language (XML) for easy parsing, which facilitates natural integration into an AWE system based on essays encoded in HTML.
@@
In §2, we describe our methods, including the data used, the models, and the approach. §The performance of the annotation models and the scoring models are presented in §3. We will discuss the findings and suggest future directions in §4.
@@ 2.1 Data
Table 1: The grade level and length statistics for the training and testing splits for the PERSUADE corpus, and the counts of essays that are responses to prompts that are dependent (Dep.) on source material and independent (Ind.) of source material.
@@ 2.1.2 Scores
Each essay was graded against standardized SAT holistic essay scoring rubric which was slightly modified for the source-based essays  [^2] https://www.kaggle.com/datasets/davidspencer/persuade-rubric-holistic-essay-scoring. Based on the rubric, a high-scoring essay (5-6) demonstrates mastery through effective development of a clear point of view, strong critical thinking with appropriate supporting evidence, well-organized structure with coherent progression of ideas, skillful language use with varied vocabulary and sentence structure, and minimal grammatical errors. In contrast, a lower-scoring essay (1-3) exhibits significant weaknesses: vague or limited viewpoint, weak critical thinking with insufficient evidence, poor organization resulting in disjointed presentation, limited vocabulary with incorrect word choices, frequent sentence structure problems, and numerous grammatical errors that interfere with meaning.
@@
The key differentiators in the rubric between high and low scoring essays are the clarity of thought, quality of supporting evidence, organizational coherence, and technical proficiency in language use. The premise behind the approach is that organizational coherence and technical proficiency in language are both made clearer by highlighting the argumentative components and conventions-based errors. Provided our pipeline for annotating essays is sufficiently accurate, these annotations should help the engine align scores in with the rubric.
@@ 2.1.3 Augmented Data
The input into the scoring model was augmented to use the annotation information using Extensible Markup Language (XML). We have an XML tag per argumentative component type. To annotate conventions errors, we used the ERRANT tool Bryant et al. (2017). The ERRANT tool classifies errors into 25 different main types, with many of these categories appearing with three different subtypes; "R" for replace, "M" for missing, and "U" for Unnecessary. For example, one category, "PUNCT", refers to a punctuation error. We can either replace, remove, or add punctuation to a sentence to make it correct, corresponding to "R:PUNCT", "U:PUNCT", and "M:PUNCT" respectively. We refer to Bryant et al. (2017) for a full explanation of the categories.
To simplify the categories for annotation purposes, we divide all possible ERRANT annotations into three labels; <Spelling>monospace-<$\verb\<Spelling\>$ >, <PunctOrth>monospace-<$\verb\<PunctOrth\>$ >, and <Grammar>monospace-<$\verb\<Grammar\>$ >. The <Spelling>monospace-<$\verb\<Spelling\>$ > label is applied to the subcategories of "SPELL", the <PunctOrth>monospace-<$\verb\<PunctOrth\>$ > is applied to the subcategories of "PUNCT" and "ORTH", while all other categories are designated labels of <Grammar>monospace-<$\verb\<Grammar\>$ >. This means that we have a total of 10 annotation labels; 7 associated with argumentative components and 3 for conventions errors. An example of the input into the model is shown in Figure 1. Since we do not have human-annotated data, the augmented data relies on the output from an annotation model and a spell-correction model. Both of which can contribute to annotations that are less accurate.
@@
Demographics: The last detail of this dataset which makes it exceptionally well-suited to the investigation of AES in an operational setting is the accompanying demographic information. This allows us to investigate any additional potential bias introduced in modeling. We measure bias by the original operational standards defined by Williamson et al. Williamson et al. (2012). While this standard is important for many reasons, there have been numerous alternative approaches to bias Ormerod et al. (2022).
@@
The population various subgroups in the train and test split, as presented in the data, have been outlined in Table 4.
@@ 2.2 Modeling details
Since the advantages of the transformer were first celebrated Vaswani et al. (2017) and BERT was trained Devlin et al. (2018), many of the state-of-the-results can be attributed to transformer-based LLMs Wang et al. (2019). For this reason, we restrict our attention to fine-tuned transformer-based LLMs, whose architectures can be described as either being encoder, decoder, or encoder-decoder models Vaswani et al. (2017). As a general rule encoder models excel in natural language inference tasks Devlin et al. (2018), decoder models excel in generative tasks Radford et al. (2018), and encoder-decoder models use excel in translation where the task benefits from representation learning Raffel et al. (2020).
@@ 2.2.1 Annotation model
The task of annotations can be framed as a token-classification task, where each token is classified into one of eight possible labels, one for each possible argumentative component in addition to one extra label for unannotated regions. This task lends itself to an encoder based model trained as a masked language model like BERT Devlin et al. (2018). Since BERT, arguably the best performing series of models are Microsofts’ DeBERTa model series He et al. (2021). The problem with these models is that many of the essays exceed the 512 token limit after tokenization. For this reason, we turn to a newly developed long context model known as Modern BERT Warner et al. (2024).
Aside from some differences in the choices of normalization layers Xiong et al. (2020), the use of gated activation functions Shazeer (2020), and more extensive pertaining, the biggest difference in the architecture is the use of Rotational Positional Embeddings (RoPE) Su et al. (2024). To understand how RoPE works, in the original implementation attention, the output of attention is given as a function of the key vectors, $k_{i}$, query vectors, $q_{j}$, and value vectors, $v_{l}$, given by
@@
where key, query, and value vectors are functions of the embedding vectors at the first attention layer. The standard construction is that these functions be affine linear functions (linear with a bias term) where the positional embedding is the addition of token embeddings and some learnable positional embedding terms. Instead of adding a vector, we define the query and key vectors using
@@
The ModernBERT model, like BERT, is trained as a masked-language model with an encoder and a token classification head. The structure of the encoder includes multiple layers of self-attention in which the attention layers alternate between rotary embeddings with base $\vartheta=1\times 10^{4}$ and $\vartheta=1.6\times 10^{5}$, however, the training was done in two phases. By adjusting the value of $\vartheta$, researchers developed a way to scale the context length of the RoPE embedding Fu et al. (2024), which has been applied to many other models AI@Meta (2024). Using this technique, modernBERT model was pretrained on 1.7 trillion tokens with a context length of 1024 with a $\vartheta=10^{-4}$, which was extended to 8196 by additional training with altered layers in which $\vartheta=1.6\times 10^{5}$. In this manner, one way of thinking of the change in $\vartheta$ values in the encoder is that the attention mechanism alternates between global and local attention.
It should be noted that many architectures have attempted to circumvent the context limitation, such as the Reformer Kitaev et al. (2020), Longformer Beltagy et al. (2020), Transformer-XL Dai et al. (2019), and XLNet Yang et al. (2019) to name a few. Many of these solutions use some sort of sliding context window and/or a recurrent adaptation of the transformer achitecture. Extending the context using rotary positional embeddings is more computationally efficient and effective at scaling to large context lengths, making Modern BERT a more appropriate choice in this context.
@@ 2.2.2 Spelling and Grammar
The most successful and accurate way to perform GEC has been to utilize representation learning, hence, we seek an encoder-decoder model, which are also sequence-to-sequence models Sutskever et al. (2014). The premise behind this method is that we are able to use the encoder to map sentences to a vector space that encodes the semantic information, while a decoder maps from the vector space to grammatically correct text Rothe et al. (2021). This suggests we use a T5 model, pretrained as a text-to-text transformer and fine-tuned to perform grammatical error correction (GEC) Martynov et al. (2023). Once the correction is defined, the original sentence and the correction are used as input into the ERRANT tool to produce a classified correction Korre and Pavlopoulos (2020).
@@ 2.2.3 Scoring Models
Given the model input includes both the essay and any annotations, we require long-context models. While we have experimented with the use of QLoRA trained generative models Ormerod and Kwako (2024), to simplify the presentation, we use the Modern BERT model for scoring in addition to annotation. In this case, the scoring models were constructed by appending a linear classification head to the ModernBERT model with 6 targets; one for each score point. We trained the models us
@@ 2.3 Evaluation
We have two models to evaluate; an annotation model and a classification model. There are many challenges to assessing annotations for argumentative clauses. We need to carefully define what it means for a particular clause to be identified and correctly classified given certain identifications may not perfectly align with the predicted components. For holistic scoring, there are much more well-defined and accepted standards presented by Williamson et al. Williamson et al. (2012).
@@ 2.3.3 Automated Scoring Evaluations
For a scoring model, the standards for agreement are specified by Williamson et al. Williamson et al. (2012). The first and primary statistic used to describe agreement is Cohen’s quadratic weighted kappa (QWK) Cohen (1960). Given scores between $1$ and $N$, we define the weighted kappa statistic by the formula
@@
$w_{ij}=\dfrac{(i-j)^{2}}{(N+1)^{2}}.$ (7)
The QWK takes values from $-1$ and $1$, indicating perfect disagreement and agreement respectively. It is often interpreted as the probability of agreement beyond random chance. The second statistic used is exact agreement, which is viewed as less reliable since it can be skewed by uneven score distributions.
The last statistic used in standardized mean difference (SMD). If $y_{t}$ represented the true score and $y_{p}$ represents the predicted score, then the SMD is given by
@@
This statistic can be interpreted as a standardized relative bias. A positive or negative value indicates that the model is introducing some positive or negative bias in the modeling process respectively. Furthermore, when we restrict this calculation to scores for a specific demographic, assuming that demographic is sufficiently well-represented, the SMD is considered a gauge of the bias associated with the modeling for that subgroup.
@@ 3.1 Annotation Accuracy
Using the 50% overlap rule, we present the number of true positives, false positives, and false negatives and the resulting F1 score for each of the component types, excluding unannotated. These results are presented in Table 5.
@@
These scores are exceptionally high with the lowest performance given by the annotators ability to discern counterclaims and rebuttals. As an indication of the annotated errors in language, the pipeline highlighted 2795 spelling errors, 1401 grammatical errors, and 201 punctuation or orthography errors. The pipeline used does not seem to be uncovering as many errors as expected.

### v2

@@ Abstract
This study illustrates how incorporating feedback-oriented annotations into the scoring pipeline can enhance the accuracy of automated essay scoring (AES). This approach is demonstrated with the Persuasive Essays for Rating, Selecting, and Understanding Argumentative and Discourse Elements (PERSUADE) corpus. We integrate two types of feedback-driven annotations: those that identify spelling and grammatical errors, and those that highlight argumentative components. To illustrate how this method could be applied in real-world scenarios, we employ two LLMs to generate annotations – a generative language model used for spell correction and an encoder-based token-classifier trained to identify and mark argumentative elements. By incorporating annotations into the scoring process, we demonstrate improvements in performance using encoder-based large language models fine-tuned as classifiers.
@@ 1 Introduction
Automated Essay Scoring (AES) uses statistical models to assign grades to essays that approximate hand-scoring Shermis and Hamner (2013). Automated Writing Evaluation (AWE) is the provision of automated feedback designed to help students iteratively improve their essays Huawei and Aryadoust (2023). Initial attempts at AES and AWE were based on Bag-of-Words (BoW) models that combine frequency-based and hand-crafted features Attali and Burstein (2006); Page (2003). Well-designed features can serve two purposes: to improve scoring accuracy and provide feedback to students to improve their essays. These features tend to be global features, such as the number of words, sentence length, or readability metrics, and are not based on fine-grained semantics or the organizational structure of essays.
Many modern AES engines employ transformer-based Large Language Models (LLM)s Rodriguez et al. (2019), which offer improved accuracy over bag-of-words models. However, this comes at the expense of reduced interpretability due to implicit feature definition. Some researchers have sought to combine LLM-derived features with traditional hand-crafted features to enhance accuracy and provide some level of interpretability Uto and Uchida (2020). LLMs also offer the ability to provide semantically rich feedback, such as key phrases from explainable AI Boulanger and Kumar (2020), annotation schemas for automated writing evaluation systems Crossley et al. (2022); Lottridge et al. (2024), and the generation of detailed feedback through LLM prompting techniques Lee et al. (2024). This study considers how these semantically rich features, designed primarily for providing feedback, can also enhance scoring accuracy.
@@
The two classes of features we consider are derived from Grammatical Error Correction (GEC) and Computational Argumentation. The goal of GEC is to provide a mapping from a sentence that may or may not contain errors in language, to a version with the same meaning with fewer language errors Martynov et al. (2023), while computational argumentation seeks to isolate and analyze the set of argumentative components of an essay Stab and Gurevych (2014a). For these features to be incorporated into an AES pipeline, we leverage the ability of language models to accurately annotate argumentative components Ormerod et al. (2023) and correct spelling and grammatical errors Rothe et al. (2021). These spelling and grammatical corrections can then be annotated and classified to provide locally defined information on conventions Korre and Pavlopoulos (2020). Once these annotations have been derived, we incorporate the annotations using Extensible Markup Language (XML) for easy parsing, which facilitates natural integration into an AWE system based on essays encoded in HTML.
@@
In §2, we describe our methods, including the data used, the models, and the approach. The performance of the annotation models and the scoring models are presented in §3. We will discuss the findings and suggest future directions in §4.
@@ 2.1 Data
Table 1: The grade level and length statistics for the training and testing splits for the PERSUADE corpus, and the counts of essays that are responses to prompts that are dependent (Dep.) on source material and independent (Ind..) of source material.
@@ 2.1.2 Scores
Each essay was graded against a standardized SAT holistic essay scoring rubric, which was slightly modified for the source-based essays  [^2] https://www.kaggle.com/datasets/davidspencer/persuade-rubric-holistic-essay-scoring. Based on the rubric, a high-scoring essay (5-6) demonstrates mastery through effective development of a clear point of view, strong critical thinking with appropriate supporting evidence, well-organized structure with coherent progression of ideas, skillful language use with varied vocabulary and sentence structure, and minimal grammatical errors. In contrast, a lower-scoring essay (1-3) exhibits significant weaknesses: vague or limited viewpoint, weak critical thinking with insufficient evidence, poor organization resulting in disjointed presentation, limited vocabulary with incorrect word choices, frequent sentence structure problems, and numerous grammatical errors that interfere with meaning.
@@
The key differentiators in the rubric between high and low scoring essays are the clarity of thought, quality of supporting evidence, organizational coherence, and technical proficiency in language use. The premise behind the approach is that organizational coherence and technical proficiency in language are both made clearer by highlighting the argumentative components and conventions-based errors. Provided our pipeline for annotating essays is sufficiently accurate, these annotations should help the engine align scores with the rubric.
@@ 2.1.3 Augmented Data
The input into the scoring model was augmented to use the annotation information using Extensible Markup Language (XML). We have an XML tag per argumentative component type. To annotate conventions errors, we used the ERRANT tool Bryant et al. (2017). The ERRANT tool classifies errors into 25 different main types, with many of these categories appearing with three different subtypes: "R" for replace, "M" for missing, and "U" for Unnecessary. For example, one category, "PUNCT", refers to a punctuation error. We can either replace, remove, or add punctuation to a sentence to make it correct, corresponding to "R:PUNCT", "U:PUNCT", and "M:PUNCT", respectively. We refer to Bryant et al. (2017) for a full explanation of the categories.
To simplify the categories for annotation purposes, we divide all possible ERRANT annotations into three labels: $\verb\<Spelling\>$, $\verb\<PunctOrth\>$, and $\verb\<Grammar\>$. The $\verb\<Spelling\>$ label is applied to the subcategories of "SPELL", the $\verb\<PunctOrth\>$ is applied to the subcategories of "PUNCT" and "ORTH", while all other categories are designated as having labels of $\verb\<Grammar\>$. This means that we have a total of 10 annotation labels: 7 associated with argumentative components and 3 for convention errors. An example of the input into the model is shown in Figure 1. Since we do not have human-annotated data, the augmented data relies on the output from an annotation model and a spell-correction model, both of which can contribute to annotations that are less accurate.
@@
Demographics: The last detail of this dataset, which makes it exceptionally well-suited to the investigation of AES in an operational setting, is the accompanying demographic information. This allows us to investigate any additional potential bias introduced in modeling. We measure bias by the original operational standards defined by Williamson et al. Williamson et al. (2012). While this standard is important for many reasons, there have been numerous alternative approaches to bias Ormerod et al. (2022).
@@
The population of various subgroups in the train and test split, as presented in the data, have been outlined in Table 4.
@@ 2.2 Modeling details
Since the advantages of the transformer were first celebrated Vaswani et al. (2017) and BERT was trained Devlin et al. (2018), many of the state-of-the-art results can be attributed to transformer-based LLMs Wang et al. (2019). For this reason, we restrict our attention to fine-tuned transformer-based LLMs, whose architectures can be described as encoder, decoder, or encoder-decoder models Vaswani et al. (2017). As a general rule, encoder models excel in natural language inference tasks Devlin et al. (2018), decoder models excel in generative tasks Radford et al. (2018), and encoder-decoder models excel in translation, where the task benefits from representation learning Raffel et al. (2020).
@@ 2.2.1 Annotation model
The task of annotations can be framed as a token-classification task, where each token is classified into one of eight possible labels, one for each possible argumentative component in addition to one extra label for unannotated regions. This task lends itself to an encoder-based model trained as a masked language model like BERT Devlin et al. (2018). Since BERT, arguably the best performing series of models are Microsofts’ DeBERTa model series He et al. (2021). The problem with these models is that many of the essays exceed the 512 token limit after tokenization. For this reason, we turn to a newly developed long context model known as ModernBERT Warner et al. (2024).
Aside from some differences in the choices of normalization layers Xiong et al. (2020), the use of gated activation functions Shazeer (2020), and more extensive pretaining, the biggest difference in the architecture is the use of Rotational Positional Embeddings (RoPE) Su et al. (2024). To understand how RoPE works, in the original implementation of attention, the output of attention is given as a function of the key vectors, $k_{i}$, query vectors, $q_{j}$, and value vectors, $v_{l}$, given by
@@
where key, query, and value vectors are functions of the embedding vectors at the first attention layer. The standard construction is that these functions be affine linear functions (linear with a bias term), where the positional embedding is the addition of token embeddings and some learnable positional embedding terms. Instead of adding a vector, we define the query and key vectors using
@@
The ModernBERT model, like BERT, is trained as a masked-language model with an encoder and a token classification head. The structure of the encoder includes multiple layers of self-attention in which the attention layers alternate between rotary embeddings with base $\vartheta=1\times 10^{4}$ and $\vartheta=1.6\times 10^{5}$. However, the training was done in two phases. By adjusting the value of $\vartheta$, researchers developed a way to scale the context length of the RoPE embedding Fu et al. (2024), which has been applied to many other models AI@Meta (2024). Using this technique, the ModernBERT model was pretrained on 1.7 trillion tokens with a context length of 1024 with a $\vartheta=10^{-4}$, which was extended to 8196 by additional training with altered layers in which $\vartheta=1.6\times 10^{5}$. In this manner, one way of thinking of the change in $\vartheta$ values in the encoder is that the attention mechanism alternates between global and local attention.
It should be noted that many architectures have attempted to circumvent the context limitation, such as the Reformer Kitaev et al. (2020), Longformer Beltagy et al. (2020), Transformer-XL Dai et al. (2019), and XLNet Yang et al. (2019), to name a few. Many of these solutions use some sort of sliding context window and/or a recurrent adaptation of the transformer architecture. Extending the context using rotary positional embeddings is more computationally efficient and effective at scaling to large context lengths, making ModernBERT a more appropriate choice in this context.
@@ 2.2.2 Spelling and Grammar
The most successful and accurate way to perform GEC has been to utilize representation learning, hence, we seek an encoder-decoder model, which is a sequence-to-sequence model Sutskever et al. (2014). The premise behind this method is that we are able to use the encoder to map sentences to a vector space that encodes the semantic information, while a decoder maps from the vector space to grammatically correct text Rothe et al. (2021). This suggests we use a T5 model, pretrained as a text-to-text transformer and fine-tuned to perform grammatical error correction (GEC) Martynov et al. (2023). Once the correction is defined, the original sentence and the correction are used as input into the ERRANT tool to produce a classified correction Korre and Pavlopoulos (2020).
@@ 2.2.3 Scoring Models
Given the model input includes both the essay and any annotations, we require long-context models. While we have experimented with the use of QLoRA-trained generative Models Ormerod and Kwako (2024), to simplify the presentation, we use the ModernBERT model for scoring in addition to annotation. In this case, the scoring models were constructed by appending a linear classification head to the ModernBERT model with 6 targets, one for each score point.
@@ 2.3 Evaluation
We have two models to evaluate: an annotation model and a classification model. There are many challenges to assessing annotations for argumentative clauses. We need to carefully define what it means for a particular clause to be identified and correctly classified, given that certain identifications may not perfectly align with the predicted components. For holistic scoring, there are many more well-defined and accepted standards presented by Williamson et al. Williamson et al. (2012).
@@ 2.3.3 Automated Scoring Evaluations
For the scoring model, we use the standards for agreement specified by Williamson et al. Williamson et al. (2012). The first and primary statistic used to describe agreement is Cohen’s quadratic weighted kappa (QWK) Cohen (1960). Given scores between $1$ and $N$, we define the weighted kappa statistic by the formula
@@
$w_{ij}=\dfrac{(i-j)^{2}}{(N-1)^{2}}.$ (7)
The QWK takes values from 1-1 and $1$, indicating perfect disagreement and agreement, respectively. It is often interpreted as the probability of agreement beyond random chance. The second statistic used is exact agreement, which is viewed as less reliable since uneven score distributions can skew it.
The last statistic used is the standardized mean difference (SMD). If $y_{t}$ represents the true score and $y_{p}$ represents the predicted score, then the SMD is given by
@@
This statistic can be interpreted as a standardized relative bias. A positive or negative value indicates that the model is introducing some positive or negative bias in the modeling process, respectively. Furthermore, when we restrict this calculation to scores for a specific demographic, assuming that demographic is sufficiently well-represented, the SMD is considered a gauge of the bias associated with the modeling for that subgroup.
@@ 3.1 Annotation Accuracy
Using the 50% overlap rule, we present the number of true positives, false positives, and false negatives, and the resulting F1 score for each of the component types, excluding unannotated. These results are presented in Table 5.
@@
These scores are exceptionally high, with the lowest performance given by the annotator’s ability to discern counterclaims and rebuttals. As an indication of the annotated errors in language, the pipeline highlighted 2795 spelling errors, 1401 grammatical errors, and 201 punctuation or orthography errors. The pipeline used does not seem to be uncovering as many errors as expected.